# Tradeoffs Between Alignment and Helpfulness in Language Models with Representation Engineering

### v3

@@ 1 Introduction
Since then, there has been an increasing body of work using this method. Wang et al. (2024b) use extracted safety vectors for inference time alignment for harmlessness, reducing jailbreaking success rate from over 30% with prompting and over 10% in supervised fine tuning to below one percent. Wang et al. (2024a) uses a method of editting model parameters that maximize the difference between toxic and untoxic responses to detoxify it. Wei et al. (2024) find sparse regions in parameter space that affect alignment brittleness, to be removed for better alignment. Marks et al. (2024) interpret causal graphs in language models and edit them to improve behaviors. van der Weij et al. (2024) extend activation steering to multiple behaviors. To improve low rank finetuning, Wu et al. (2024) utilize a procedure of tuning representations directly to substantially reduce the trainable parameters of finetuning compared to LoRA. Xu et al. (2024); Li et al. (2024) use concept activation vectors to jailbreak, they also observe that concepts that activate different behaviors are linearly separable. Zhang et al. (2024) remove hallucinations by editting truthfulness concepts.
@@
Understanding the tradeoff between model helpfulness and alignment is important for designing safe yet useful LLM systems. In this work we aim to shed light on the benefits and limitations of representation engineering for LLM alignment, i.e., how much does alignment improve with this method and what is the cost in terms of the model’s abilities. We approach this question theoretically at first, and then provide empirical evidence for the validity of our theory.
@@
In sections 2 and 3, we set up our theoretical framework and present our theoretical results respectively. We find that representation engineering increases alignment linearly with the representation engineering vector norm (theorem 1), while the helpfulness of the model, defined as the probability of answering a query correctly, decreases quadratically with the representation engineering vector norm (theorem 2). Consequently, alignment can be guaranteed with large enough representation injections (corollary 1), though at the cost of significantly reducing the model’s helpfulness, e.g.multiple choice question answering reduced to random guessing (corollary 3). Conversely, this means that when injecting vectors of small norms, the improvement of alignment is initially faster than the decrease in helpfulness, possibly indicating that there is a regime where prompt engineering is more effective, allowing for inference time alignment while maintaining the model’s helpful capabilities. See figure 1 for an illustration of this intuition.
In section 4 we explore the validity of our assumptions and results in an experimental setting: We calculate alignment, as defined by the theoretical framework, as a function of representation engineered vector norms corresponding to the desired behaviors and find that it increases as predicted by theorem 1. This is done by aligning with representation engineering an unaligned (pretrained) model with respect to desired behaviors (“harmless", “not-racist"), and misaligning an aligned (RLHF) model to undesired behaviors (“harmful", “racist"). Then, we calculate the helpfulness of the model, quantified by its question answering abilities over different knowledge domains, with the same aligning vectors, and find that the decay with increased vector norm described in theorem 2 is manifested. Furthermore, injecting large norms of these vectors leads to significant reduction of helpfulness, such as performance on multiple-choice questions that is equivalent to random guessing (corollary 3). Together, these results correspond to the intuitive illustration in figure 1 .
@@ 2 Preliminaries
Parameterizing representation engineering is done by modifying each layer’s hidden state via adding a corresponding engineered vector: Denote a representation engineered model $\mathbb{P}_{\theta}$ of $L$ layers, with a set of engineered representations $R_{e}=(r_{e}^{(l=1)},...,r_{e}^{(l=L-1)})$ by $\mathbb{P}_{\theta,r_{e}}$, which is applied by adding to each hidden state the corresponding engineered vector:
@@
To quantify alignment, we use the behavior expectation definition of alignment as in Wolf et al. (2023). We will use a binary behavior scoring function:
@@
Let $B:\Sigma^{*}\rightarrow\{-1,+1\}$ be a binary behavior scoring function, the behavior of a prompted model $\mathbb{P}(\cdot|q)$ is defined as:
$B[\mathbb{P}(\cdot|q)]=\mathbb{E}_{a\sim\mathbb{P}(\cdot|q)}[B(a)]=\sum_{a_{+}\in aligned}\mathbb{P}_{\theta}(a_{+}|q)-\sum_{a_{-}\in misaligned}\mathbb{P}_{\theta}(a_{-}|q)$ (3)
Notice that while $B$ is a binary function, the behavior expectation is in the range $[-1,+1]$. A binary behavior scoring function reflects cases where an aligned response is required and unaligned responses must be filtered. In theorem 1 we will prove that representation engineering is an effective method by lower bounding the behavior expectation. Notice that high probability of outputting a positive/negative response gives a positive/negative contribution to the behavior expectation, thus the sign and absolute value of behavior expectation is a good measure for the alignment of a model.
We define the model’s helpfulness as its ability to produce correct answers for queries. This can be measured as the likelihood of outputing a correct answer to a query. In order to test helpfulness for a specific skill, we can look at the probabiltiy for a specific query. See figure 1 for an example. The helpfulness of a model with respect to a specific query is defined as:
@@
Where $P_{\theta}(a_{correct}|q)$ is the model’s probability of outputting the correct answer $a$ to the query $q$. By this definition, the helpfulness is in the range $[0,1]$. The motivation behind this definition is to quantify the general capabilities of the model when engineered representations are injected into it.
Ideally, a model that interacts with a user should be both aligned and helpful, meaning its response is appropriate w.r.t. a desired behavior (i.e., positive behavior expectation) and also useful (i.e., high probability of giving a correct answer to the query).
Lastly, note the use of the model’s next token prediction to determine alignment and helpfulness. Initially this may seem to restrict the generality of our results, but note that for alignment, it is often the first token that determines whether an answer is aligned or not ("Sorry, as an AI I cannot…" vs "Sure! Here’s…"). For helpfulness this point is more subtle, as some tasks are more intricate, however, as demonstrated by our experiments, there are numerous tasks that can be reduced to next token prediction.
@@ 3 Main Results
We will show that representation engineering improves alignment, but harms helpfulness. Theorem 1 shows that behavior expectation is bounded from below by a hyperbolic tangent function, such that it approaches $+1$ for increasing size of injected vectors and increases linearly within a bounded range. This in principle allows to sample an aligned response for any adversarial attack (corollary 1), demonstrating the power of representation engineering as an alignment technique. Theorem 2 shows that the helpfulness is maximized in the vicinity of norm zero injected vectors (i.e., no representation engineering) and in corollary 3 that as the norm is increased, the helpfulness decays to random guessing. The assumptions used to prove the theorems are presented formally in appendix A.
The following statement quantifies how alignment is improved by representation engineering. It assumes the change in all layers amounts to a good classifier of positive and negative behavior tokens in the last layer w.r.t.the query, as depicted in figure 2a. This condition was chosen due to the popular choice in representation engineering to use injected representations $\{r_{e}^{(l)}\}$, that are themselves classifiers for positive and negative representations on the intermediate layers. This is because they are learned from contrasting positive and negative behavior representations for different queries, such as mean centering, $r_{e}^{(l)}=\mathbb{E}_{good,bad}[r_{good}^{(l)}-r_{bad}^{(l)}]$, or PCA, $r_{e}^{(l)}=\text{argmax}_{v:||v||=1}[\mathbb{E}_{good,bad}|\langle v,r_{good}^{(l)}-r_{bad}^{(l)}\rangle|^{2}]$ such that they form linear classifiers for the intermediate layers due to the positive/negative inner product with positive/negative answer representations. Notably, in Xu et al. (2024) it is shown empirically that such concept classes in latent space are linearly separable. We discuss this assumption further in A and provide empirical evidence. Furthermore, the classification condition can be softened to an imperfect classifier, as discussed in appendix A and shown in appendix in F, to yield similar results.
@@
Let $\mathbb{P}_{\theta,r_{e}}(\cdot|q)$ be a model prompted with query $q$ and injected with representations of coefficient $r_{e}$. Let $B:\Sigma^{*}\rightarrow\{-1,+1\}$ be a behavior scoring function. The injections to all layers amounts to a change in the final hidden layer representation that is $q$ dependent, denoted by the vector $\delta r^{(L)}_{e}(q)$. Let $\Delta$ be the margin of classification over the model’s representations for aligned and misaligned answers with respect to $B$, when using $\delta r^{(L)}_{e}(q)$ as a classifier. Then, the behavior expectation of the model conditioned on the query $q$ satisfies:
$B[\mathbb{P}_{\theta,r_{e}}(\cdot|q)]\geq tanh(\Delta\lambda\cdot r_{e}+arctanh(B_{0}))$ (5)
Where $B_{0}=B[\mathbb{P}_{\theta}(\cdot|q)]$ is the behavior expectation without representation engineering and $\lambda$ is a model dependent coefficient relating between $r_{e}$ and the corresponding final hidden state norm.
As can be seen in the mathematical expression and in figure 2b for $B_{0}=-0.5$, this lower bound is a shifted hyperbolic tangent function w.r.t $r_{e}$. At $r_{e}=0$ the bound gives $B_{0}$, which is the unaltered model’s behavior. As $r_{e}$ is increased, the bound approaches $+1$, meaning the behavior asymptotically approaches $+1$. We see that for $B_{0}$ that is not too close $-1$, the increase in behavior expectation is linear due to the hyperbolic tangent’s nature, while if it is very close to $-1$, $r_{e}$ is to be increased before seeing the linear effect. Thus for behaviors on which the model is negative but also has a small tendency for positive answers, the linear effect should be felt near $r_{e}=0$. In section 4, we present our numerical estimation $\Delta\lambda$ in the range $0.1-3$, both based on the linear classifier condition and direct alignment measurement. For proof see appendix section B.
@@
We note that in contrast to Wolf et al. (2023), that has a framework centralized on using prompts to misalign frozen models, i.e.models whose weights and representations are not changed after training, here the model is not frozen due to representation engineering, and accordingly a different result is obtained on guaranteeing an aligned response - for any adversarial attack, it is possible to use large enough norms with representation engineering to produce an aligned response if the learned injected representations accumulate to a good classifier of positive and negative answer representations in the final layer.
@@
Let $\epsilon>0$, $\mathbb{P}_{\theta}$ a language model and $q$ a prompt that induces negative behavior $B[\mathbb{P}_{\theta}(\cdot|q)]<\gamma<0$ without representation engineering. Under the conditions of theorem 1, using an injected vector norm of $r_{e}>\frac{1}{\Delta\lambda}(arctanh(1-\epsilon)-arctanh(\gamma))$ leads to behavior expectation $B[\mathbb{P}_{\theta,r_{e}}(\cdot|q)]>1-\epsilon$.
Now, we shall bound from above the helpfulness of the model as a function of representation engineering. Intuitively, editting the model’s representation in a specific direction adds random noise to other latent concepts of the model, causing a degradation in its other capabilities. The following theorem formally states this.
@@
Let $\mathbb{P}_{\theta,r_{e}}(\cdot|q)$ be a model prompted with query $q$ and injected with representations of coefficient $r_{e}$. If the directionality of the last hidden layer representation change due to the injections in all layers, distributes normally with variance $\sigma^{2}$ w.r.t. the correct and incorrect answer representations, the helpfulness of the model on the query is bounded with probability $1-\frac{2}{|V|}$ by:
$\mathbb{P}_{\theta,r_{e}}(a_{correct}|q)\leq\frac{P_{0}}{P_{0}+(1-P_{0})\cdot\alpha(1+\frac{\lambda^{2}\sigma^{2}\beta^{2}}{2}r_{e}^{2})}$ (6)
Where $P_{0}=\mathbb{P}_{\theta,r_{e}=0}(\cdot|q)$ is the probability of answering correctly without representation engineering, $|V|$ is the vocabulary size and $\alpha,\beta>0$ that depend on the query. $\lambda$ is a model dependent coefficient relating between $r_{e}$ and the corresponding final hidden state norm.
The proof is presented in appendix C. The above bound is illustrated in figure 3 for different values of $\beta$. As can be seen, around $r_{e}=0$, the bound is parabolic, i.e.  the decrease is proportional to $-r_{e}^{2}$, this can be obtained by expanding the bound near $r_{e}=0$. On the other hand, for large $r_{e}$, we see a decay to zero at a rate proportional to $r_{e}^{-2}$, this can be obtained by expanding the bound for large $r_{e}$.
@@
The parameter $\alpha\in[0,1]$ measures the tightness of the bound at $r_{e}=0$, since the true helpfulness at $r_{e}=0$ is $P_{0}$, while our helpfulness bound is $\frac{P_{0}}{P_{0}+\alpha(1-P_{0})}$. Thus $\alpha=1$ means the bound at $r_{e}=0$ coincides with the true helpfulness, while smaller $\alpha$ means the bound overshoots the true helpfulness. In our results, we obtain $\alpha\leq 0.5$. Figure 3 depicts this overshooting for $\alpha=0.25$. Even so, as explained above, the tightness is at least $1-P_{0}$ regardless of $\alpha$, so it is always meaningful for queries the model is initially helpful on.
The product of parameters $\lambda\sigma\beta$ are a measure for the rate/curvature of the quadratic decay, as they are the coefficient multiplying $r_{e}^{2}$. $\lambda$ is the same scaling parameter from theorem 1, $\sigma$ represents the standard deviation of random noise added to the logits due to representation engineering (depicted in figure 3a and formally defined in A), and $\beta$ is the minimum between two weighted sums of half normal distributions with parameter $\sigma^{\prime}=1$. In section 4, we present our empirical estimation $\lambda\sigma\beta$ in the range $0.1-0.66$, both based on the normal logit noise condition and direct helpfulness measurement. Hence the decay should be felt at coefficients $r_{e}$ of order of size $1$.
Figure 3: (a) Directionality of change to last hidden layer due to representation engineering distributes normally with variance $\sigma^{2}$ w.r.t. correct and incorrect answer representations. (b) Plot of helpfulness bound with given parameters of $P_{0}$, $\alpha$ and $\lambda\sigma\beta$.
From this, we can derive a bound on the expectation value of helpfulness on a dataset:
Corollary 2.
Under the conditions of theorem 2, the expectation value of helpfulness on a dataset is bounded by:
$\mathbb{E}_{q\in dataset}[\mathbb{P}_{\theta,r_{e}}(a_{correct}|q)]\leq(1-\frac{1}{|V|})\frac{1}{1+\alpha^{\prime}(1+\frac{\sigma^{2}\lambda^{2}\beta^{\prime 2}}{2}r_{e}^{2})}+\frac{1}{|V|}$ (7)
For $\alpha^{\prime},\beta^{\prime}>0$ that depend on the dataset.
So, for datasets that have high helpfulness expectation value (about $0.9$ or higher), this bound shows helpfulness decay quadratically.
@@
Under the conditions of theorem 2, the expected value of the helpfulness on a dataset of queries, $\mathbb{E}_{q\in dataset}[\mathbb{P}_{\theta,r_{e}}(a_{correct}|q)]$ is asymptotically bounded from above by $\frac{1}{N}$ as $|r_{e}|\rightarrow\infty$. Where $N$ is the number of possible answers for each query.

### v4

@@ 1 Introduction
Since then, there has been an increasing body of work using this method. Wang et al. (2024b) use extracted safety vectors for inference time alignment for harmlessness, reducing jailbreaking success rate from over 30% with prompting and over 10% in supervised fine tuning to below one percent. Wang et al. (2024a) uses a method of editing model parameters that maximize the difference between toxic and untoxic responses to detoxify it. Wei et al. (2024) find sparse regions in parameter space that affect alignment brittleness, to be removed for better alignment. Marks et al. (2024) interpret causal graphs in language models and edit them to improve behaviors. van der Weij et al. (2024) extend activation steering to multiple behaviors. To improve low rank finetuning, Wu et al. (2024) utilize a procedure of tuning representations directly to substantially reduce the trainable parameters of finetuning compared to LoRA. Xu et al. (2024); Li et al. (2024) use concept activation vectors to jailbreak, they also observe that concepts that activate different behaviors are linearly separable. Zhang et al. (2024) remove hallucinations by editing truthfulness concepts. Additionally, the method scales to SOTA models, such as Sonnet’s Claude Templeton (2024), using a similar method of sparse auto encoders, which extracts interpretable features from the model that can be used to manipulate the model through steering.
@@
Understanding the tradeoff between model helpfulness and alignment is important for designing safe yet useful LLM systems. Previous empirical works have shown tradeoffs between quality and diversity and between helpfulness and safety in LLMs due to instruct finetuning (Florian et al., 2024; Bianchi et al., 2023; Röttger et al., 2023), and reduction in performance due to watermarking (Ajith et al., 2023). In this work we aim to shed light on the benefits and limitations of representation engineering for LLM alignment, i.e., how much does alignment improve with this method and what is the cost in terms of the model’s abilities. We approach this question theoretically at first, and then provide empirical evidence for the validity of our theory.
@@
In sections 2 and 3, we set up our theoretical framework and present our theoretical results respectively. We find that representation engineering increases alignment linearly with the representation engineering vector norm (theorem 1), while the helpfulness of the model, defined as the probability of answering a query correctly, decreases quadratically with the representation engineering vector norm (theorem 2). Consequently, alignment can be guaranteed with large enough representation injections (corollary 1), though at the cost of significantly reducing the model’s helpfulness, e.g.multiple choice question answering reduced to random guessing (corollary 2). Conversely, this means that when injecting vectors of small norms, the improvement of alignment is initially faster than the decrease in helpfulness, possibly indicating that there is a regime where representation engineering is more effective, allowing for inference time alignment while maintaining the model’s helpful capabilities. See figure 1 for an illustration of this intuition.
In section 4 we explore the validity of our assumptions and results in an experimental setting: We calculate alignment, as defined by the theoretical framework, as a function of representation engineered vector norms corresponding to the desired behaviors and find that it increases as predicted by theorem 1. This is done by aligning with representation engineering an unaligned (pretrained) model with respect to desired behaviors (“harmless", “not-racist"), and misaligning an aligned (RLHF) model to undesired behaviors (“harmful", “racist"). Then, we calculate the helpfulness of the model, quantified by its question answering abilities over different knowledge domains and coding capabilities, with the same aligning vectors, and find that the decay with increased vector norm described in theorem 2 is manifested. Furthermore, injecting large norms of these vectors leads to significant reduction of helpfulness, such as performance on multiple-choice questions that is equivalent to random guessing (corollary 2). Together, the results correspond to the intuitive illustration in fig. 1.
@@ 2 Preliminaries
Parameterizing representation engineering is done by modifying each layer’s hidden state via adding a corresponding engineered vector: Denote a representation engineered model $P_{\theta}$ of $L$ layers, with a set of engineered representations $R_{e}=(r_{e}^{(l=1)},...,r_{e}^{(l=L-1)})$ by $P_{\theta,r_{e}}$, which is applied by adding to each hidden state the corresponding engineered vector:
@@
To quantify alignment, we use the behavior expectation definition of alignment as in Wolf et al. (2023). We will use a binary scoring function, with labels ±$\pm 1$ for aligned/misaligned answers. The results can be extended to more complex behavior scoring functions over $[-1,+1]$, to yield qualitatively similar results, as discussed appendix H:
@@
Let $B:\Sigma^{*}\rightarrow\{-1,+1\}$ be a binary behavior scoring function, the behavior of a prompted model $P(\cdot|q)$ is defined as:
$B[P_{\theta}(\cdot|q)]=\mathbb{E}_{a\sim P_{\theta}(\cdot|q)}[B(a)]=\sum_{a_{+}\in aligned}P_{\theta}(a_{+}|q)-\sum_{a_{-}\in misaligned}P_{\theta}(a_{-}|q)$ (3)
Notice that while $B$ is a binary function, the behavior expectation is in the range $[-1,+1]$, reflecting cases where an aligned response is required and unaligned responses must be filtered. In theorem 1 we will prove that representation engineering is an effective alignment method by lower bounding the behavior expectation. Notice that high probability of outputting a positive/negative response gives a positive/negative contribution to the behavior expectation, thus the sign and absolute value of behavior expectation is a good measure for the alignment of a model.
The model’s helpfulness can be quantified as its ability to produce useful answers to user’s queries (knowledge questions, code generation, summarization, etc.). In order to theoretically analyze helpfulness, we focus on queries where correctness can be defined, such as knowledge based question answering on various domains (see figure 1 for an example), or producing code to solve a problem. This can be measured as the likelihood of outputting a correct answer to a query:
@@
Where $P_{\theta}(a_{correct}|q)$ is the model’s probability of outputting the correct answer $a$ to the query $q$. By this definition, the helpfulness is in the range $[0,1]$, and motivation behind it is to quantify the general capabilities of the model when engineered representations are injected into it. For queries where correctness is not defined, the bounds we derive are expected to still be meaningful as they can also describe the rate of the model’s deviation from its original distribution due to representation engineering, as will be explained in the next section.
Ideally, a model that interacts with a user should be both aligned and helpful, meaning its response is appropriate w.r.t. a desired behavior (i.e., positive behavior expectation) and also useful (i.e., high probability of giving a correct answer to the query). In the next section, we will provide results on alignment and helpfulness under the use of representation engineering, based on the model’s next token prediction, which provides simple analytical forms for alignment and helpfulness. In appendix I, we extend the results for multi-token answers, which yields qualitatively similar results, with somewhat more complex form.
@@ 3 Main Results
We will show that representation engineering improves alignment, but harms helpfulness. Theorem 1 shows that behavior expectation is bounded from below by a hyperbolic tangent function, such that it approaches $+1$ for increasing size of injected vectors and increases linearly within a bounded range. This in principle allows to sample an aligned response for any adversarial attack (corollary 1), demonstrating the power of representation engineering as an alignment technique. Theorem 2 shows that the helpfulness is maximized in the vicinity of norm zero injected vectors (i.e., no representation engineering) and in corollary 2 that as the norm is increased, helpfulness decays to random guessing. The assumptions used to prove the theorems are presented formally in appendix A.
The following statement quantifies how alignment is improved by representation engineering. It assumes the injected representations in all layers accumulate to a change in the last hidden layer representation that classifies positive and negative behavior answers to the query, as depicted in figure 2a. This condition was chosen due to the popular choice in representation engineering to use injected representations $\{r_{e}^{(l)}\}$, that are themselves classifiers for positive and negative representations on the intermediate layers. This is because they are learned from contrasting positive and negative behavior representations for different queries, such as mean centering, $r_{e}^{(l)}=\mathbb{E}_{good,bad}[r_{good}^{(l)}-r_{bad}^{(l)}]$ (Jorgensen et al. (2023)), or PCA, $r_{e}^{(l)}=\text{argmax}_{v:||v||=1}[\mathbb{E}_{good,bad}|\langle v,r_{good}^{(l)}-r_{bad}^{(l)}\rangle|^{2}]$ (Zou et al. (2023a)), such that they form linear classifiers for the intermediate layers due to the positive/negative inner product with positive/negative answer representations. Notably, in Xu et al. (2024) it is shown empirically that such concept classes in latent space are linearly separable. We discuss this assumption further in A and provide empirical evidence. Furthermore, the classification condition can be softened to an imperfect classifier, as discussed in appendix A and shown in appendix in F, to yield similar results.
@@
Let $P_{\theta,r_{e}}(\cdot|q)$ be a model prompted with query $q$ and injected with representations of coefficient $r_{e}$. Let $B:\Sigma^{*}\rightarrow\{-1,+1\}$ be a behavior scoring function. The injections to all layers amounts to a change in the final hidden layer representation that is $q$ dependent, denoted by the vector $\delta r^{(L)}_{e}(q)$. Assume that the representations of aligned and misaligned answers w.r.t. $B$ are linearly separable, and that $\delta r^{(L)}_{e}(q)$ linearly classifies them with margin $\Delta$. Then, the behavior expectation of the model conditioned on the query $q$ satisfies:
$B[P_{\theta,r_{e}}(\cdot|q)]\geq tanh(\Delta\lambda\cdot r_{e}+arctanh(B_{0}))$ (5)
Where $B_{0}=B[P_{\theta}(\cdot|q)]$ is the behavior expectation without representation engineering and $\lambda$ is a model dependent coefficient relating between $r_{e}$ and the corresponding final hidden state norm.
As can be seen in the mathematical expression and in figure 2b for $B_{0}=-0.5$, this lower bound is a shifted hyperbolic tangent function w.r.t $r_{e}$. At $r_{e}=0$ the bound gives $B_{0}$, which is the unaltered model’s behavior. As $r_{e}$ is increased, the bound approaches $+1$, meaning the behavior asymptotically approaches $+1$. We see that for $B_{0}$ that is not too close to $-1$, the increase in behavior expectation is linear due to the hyperbolic tangent’s nature, while if it is very close to $-1$, $r_{e}$ is to be increased before seeing the linear effect. Thus for behaviors on which the model is negative but also has a small tendency for positive answers, the linear effect should be felt near $r_{e}=0$. In section 4, we present our numerical estimation $\Delta\lambda$ in the range $0.1-3$, both based on the linear classifier condition and direct alignment measurement. For proof see appendix section B.
@@
In contrast to Wolf et al. (2023), that has a framework centralized on using prompts to misalign frozen models, i.e. models whose weights and representations are not changed after training, here the model is not frozen due to representation engineering, and accordingly a different result is obtained on guaranteeing an aligned response – for any adversarial attack, using large enough norms with representation engineering produces an aligned response if the learned injected representations accumulate to a good classifier of positive and negative answer representations in the final layer.
@@
Let $\epsilon>0$, $P_{\theta}$ a language model and $q$ a prompt that induces negative behavior $B[P_{\theta}(\cdot|q)]<\gamma<0$ without representation engineering. Under the conditions of theorem 1, using an injected vector norm of $r_{e}>\frac{1}{\Delta\lambda}(arctanh(1-\epsilon)-arctanh(\gamma))$ leads to behavior expectation $B[P_{\theta,r_{e}}(\cdot|q)]>1-\epsilon$.
This can be extended to multi-token answers, by enforcing the above result on each decoding step of the generated answer, as explained in appendix I. The binary behavior score can also be extended beyond binary, as explained in appendix H.
Now, we shall bound from above the helpfulness of the model as a function of representation engineering. We formally bound the probability of producing correct answers to queries where correctness is well-defined. Yet, even for queries where this is not the case, the bound can still be relevant, as it quantifies the model’s deviation from its original distribution due to representation engineering. Hence if the model was initially helpful on a task, a random deviation to its probability distribution is expected to decrease model performance proportionally to the size of the deviation.
Intuitively, editing the model’s representation in a specific direction adds random noise to other latent concepts of the model, causing a degradation in its other capabilities. This is introduced in our framework through the resulting change to the final hidden layer $\delta r_{e}(q)=r_{\theta,r_{e}}^{(L)}-r_{\theta}^{(L)}$, we will assume its direction $\frac{\delta r_{e}(q)}{|\delta r_{e}(q)|}$ contains random projections w.r.t. latent representations of correct and incorrect answers, which creates noise in the model’s distribution. The noise is expected to be random on the highest probability tokens, since they answer a query that is unrelated to the behavior being enhanced (intuitively depicted in figure 3a). We verify this empirically in appendix A.3. Thus, we assume random noise on the top $T$ tokens making up a large probability mass of the answer distribution, $1-\epsilon$, (e.g. $T\sim 10$ typically makes $\epsilon\sim 0.1$), and do not make assumptions on the rest of the vocabulary. The following theorem formally states this.
@@
Let $P_{\theta,r_{e}}(\cdot|q)$ be a model prompted with query $q$ and injected with representations of coefficient $r_{e}$. If the resulting change to the directionality of the last hidden layer representation due to the injections in all layers, distributes randomly with variance $\sigma^{2}>0$ w.r.t. the representations of correct and incorrect answers making up $1-\epsilon$ of the probability mass, the helpfulness of the model on the query is bounded with probability $1-\frac{2}{T}$ by:
$P_{\theta,r_{e}}(a_{correct}|q)\leq\frac{P_{0}}{P_{0}+(1-P_{0})\cdot\alpha(1-\epsilon)(1+\frac{\lambda^{2}\sigma^{2}\beta^{2}}{2}r_{e}^{2})}$ (6)
Where $P_{0}=P_{\theta,r_{e}=0}(\cdot|q)$ is the probability of answering correctly without representation engineering, $T$ is the number of tokens making $1-\epsilon$ of the probability mass and $\alpha,\beta>0$ that depend on the query. $\lambda$ is a model dependent coefficient relating between $r_{e}$ and the corresponding final hidden state norm.
The proof is presented in appendix C and the assumption formally defined in appendix A. The above bound is illustrated in figure 3b for different values of $\beta$. As can be seen, around $r_{e}=0$, the bound is parabolic, i.e.  the decrease is proportional to $-r_{e}^{2}$, this can be obtained by expanding the bound near $r_{e}=0$. On the other hand, for large $r_{e}$, we see a decay to zero at a rate proportional to $r_{e}^{-2}$, this can be obtained by expanding the bound for large $r_{e}$. This can be extended to multi-token answers, by enforcing the above result on each decoding step of the generated answer, as explained in appendix I.
@@
The parameter $\alpha\in[0,1]$ measures the tightness of the bound at $r_{e}=0$, since the true helpfulness at $r_{e}=0$ is $P_{0}$, while our helpfulness bound is $\frac{P_{0}}{P_{0}+\alpha(1-P_{0})}$. Thus $\alpha=1$ (and $\epsilon=0$) means the bound at $r_{e}=0$ coincides with the true helpfulness, while smaller $\alpha$ means the bound overshoots the true helpfulness. In our results, we obtain $\alpha\leq 0.5$. Figure 3 depicts this overshooting for $\alpha=0.25$. Even so, as explained above, the tightness is at least $1-P_{0}$ regardless of $\alpha$, so it is always meaningful for queries the model is initially helpful on.
The product of parameters $\lambda\sigma\beta$ are a measure for the rate/curvature of the quadratic decay, as they are the coefficient multiplying $r_{e}^{2}$. $\lambda$ is the same scaling parameter from theorem 1, $\sigma$ represents the standard deviation of random noise added to the logits due to representation engineering (depicted in figure 3a and formally defined in A), and $\beta$ is the minimum between two weighted sums of positive variables with parameter $\sigma^{\prime}=1$. In section 4, we present our empirical estimation $\lambda\sigma\beta$ in the range $0.1-0.66$, both based on the logit noise condition and direct helpfulness measurement. Hence the decay should be felt at coefficients $r_{e}$ of order of size $1$.
Figure 3: (a) Directionality of change to last hidden layer due to representation engineering distributes randomly with variance $\sigma^{2}$ w.r.t. correct and incorrect answer representations. (b) Plot of helpfulness bound with given parameters of $P_{0}$, $\alpha$ and $\lambda\sigma\beta$.
@@
Under the conditions of theorem 2, the expected value of the helpfulness on a dataset of queries, $\mathbb{E}_{q\in dataset}[P_{\theta,r_{e}}(a_{correct}|q)]$ is asymptotically bounded from above by $\frac{1}{N}$ as $|r_{e}|\rightarrow\infty$. Where $N$ is the number of possible answers for each query.